OpenStudy (anonymous):

write the expression as a complex number in standard form. 3-(2+2i)+4i

OpenStudy (mathstudent55):

To remove parentheses, you must distribute the negative outside the parentheses: 3 - (2 + 2i) + 4i = = 3 - 2 - 2i + 4i Now combine like terms

OpenStudy (anonymous):

is it going to be 5+6i

OpenStudy (mathstudent55):

What is 3 - 2 ?

OpenStudy (anonymous):

oh 1

OpenStudy (mathstudent55):

What is -2i + 4i?

OpenStudy (anonymous):

-2i

OpenStudy (mathstudent55):

Right, so the answer is 1 - 2i

OpenStudy (anonymous):

how about this one 6i+(5-3i)-5

OpenStudy (mathstudent55):

Here, ther is a plus sign to the left of the parentheses, so you cna simply remove the parentheses and combine like terms. 6i + (5 - 3i) - 5 = = 6i + 5 -3i - 5 Now combine like terms

OpenStudy (anonymous):

so it would be just -3i

OpenStudy (mathstudent55):

Like terms are terms that have the same variable part. 6i and -3i have the same variable parts, i. They can be combined together. 6i - 3i = 3i 5 and -5 have the same variable part. Both have no variable at all. 5 and -5 can be combined together. 5 - 5 = 0 So you end up with just 3i 3i, not - 3i, since it's 6i - 3i, not 3i - 6i.

OpenStudy (anonymous):

so i was almost right

OpenStudy (mathstudent55):

Yes, but almost is still not right. You need to be careful adding positive and negative numbers together.

OpenStudy (anonymous):

so if its a problem where one regular number is in the same parentheses as i number with i what do you do